Based on the table with ionization energies (eV) for 1s and 2s electrons given at the end, determine the Zeffthat is experienced by the electron in the 2s atomic orbital in Li and Be. Also, comment on your findings. H H e Li 13.6 24.6 5.4 9.3 8.311.314.5 13.6 17.4 21.6 5.1 54.4 75.6 18.2 25.2 24.429.6 35.1 35.0 41.0 47.3 122.5 153.9 37.947

The solution is 1.26 (units not mentioned) for Li and 1.65 (units not mentioned) for Be.

Hint: Substitute Zeff in Z for H-like atoms .
Answer #1

oT 久 6. RH ( ) where Rad beata t era conatant OuT ayuant um mwn ber mamt ummumFom Li (65x 10 19 8, 652x10 クスeffRH = 2-18x 10-18丁 2 r2 I g 1.49 x 10 x 4 2-18 X 10 3 2 굿eff 2 굿eff= 1.65 ラヌeff =2-734